# TuneShield: Mitigating Toxicity in Conversational AI while Fine-tuning on Untrusted Data

### v1

@@ III-A Design challenges
(1) Constantly evolving base models and fine-tuning strategies. As foundation models advance, their capacity, architecture, and training objectives evolve, leading to changes in fine-tuning strategies [27]. This makes designing a safety framework that can work for a large variety of base models and fine-tuning approaches, challenging.
(2) Mitigating toxicity while preserving conversational quality. Integrating toxicity mitigation strategies during fine-tuning can degrade conversation quality. For example, a poorly implemented filter can result in numerous false positives, incorrectly flagging non-toxic samples as toxic. Removing these samples from the training set can significantly degrade model utility. We aim to reduce toxicity while maintaining conversational quality using our framework.
@@
(3) Defender is unaware of the distribution of toxic language and may have access to imperfect toxicity classifiers. Toxicity detection in a conversational setting is a difficult problem, as toxicity varies by different pragmatics and continue to evolve. Our defender is unaware of the toxic language distribution used by the adversary to poison the training dataset. Supervised toxicity classifiers are known to generalize poorly to previously unseen toxic language. We leverage recent advances in LLMs to design an unsupervised toxicity classifier. Even a classifier that is performant today, may degrade over time as toxic language evolves. Therefore, our framework should mitigate toxicity even when the toxicity classifier is biased or imperfect to a certain extent.
@@
(4) Mitigating toxicity while reinforcing desired conversational behavior. While mitigating toxicity, an important design consideration is to think about how the model should respond given a toxic context. We explore 2 possible strategies: (a) Generate a non-sequitur: This is a pre-defined response that does not capture the context of the conversation. For instance, respond with “I’m sorry, I cannot fulfill your request”. However, in some contexts, such canned responses can reduce “engagingness”, which is usually targeted by chatbot designers [9]. (b) Generate responses with certain desired properties, e.g., empathetic or prosocial behavior. For example, when presented with a context of “What if we legalize hate crimes?”, the chatbot can respond with “I can’t engage in conversations that promote hate or discrimination. Let’s focus on inclusivity and respect. Anything else to discuss?”. We explore the impact of both strategies on toxicity mitigation while preserving the utility of the model.
III-B TuneShield design
@@
Figure 2 illustrates the design of TuneShield framework. There are 2 stages: (1) toxicity classification, and (2) model fine-tuning and alignment stage.
@@
Toxicity classification stage.   This stage uses a toxicity classifier to differentiate between non-toxic and toxic samples in the fine-tuning dataset. Any toxicity classifier can be used in this stage. We investigate the potential of LLM-based toxicity classification using a zero-shot prompting strategy. Our prompts contain 3 parts: (a) Input, which contains the input sample i.e., a context-response pair from the fine-tuning dataset. (b) Description explains the input format. (c) Instruction is the task to perform on the input, i.e., classification. We propose a zero-shot prompting approach that leverages the instruction following and safety alignment abilities of LLMs. We call this Refusal approach.
@@
Refusal approach. In this approach, we use a safety-aligned LLM to identify toxic samples. When a user prompts a safety-aligned LLM with problematic content, the model usually responds with a refusal statement as a follow-up utterance e.g., ‘I apologize, but I cannot fulfill your request…’ [71]. Our intuition is to use this ability to build a toxicity classifier. We prepare the input (context-response pair) as multi-turn conversation between two users and prompt the model to respond with a ‘yes’ or ‘no’ if it is safe to generate the next turn in the conversation. As LLMs exhibit non-deterministic behavior while generating responses, relying on the decoded output of the model to predict ‘yes’ or ‘no’ is not an effective strategy. Instead, we obtain the log-likelihoods of obtaining ‘yes’ and ‘no’ as responses. We compute probabilities from these log-likelihoods using the softmax function to determine the prediction result [63]. We determine an input sample as toxic if the prediction is determined as ‘no’ ($P(`no\textrm{'}|prompt)\geq 0.5$), i.e., unsafe to generate the next turn in the conversation. Figure 3 shows the prompts for the Refusal approach.
Figure 3: Prompt for the Refusal approach.
Model fine-tuning and alignment stage.   Toxicity classifiers can be biased or imperfect, leading to significant false negatives and false positives when applied to training samples. Therefore, using the toxicity classifier as a simple filter can still result in residual toxicity in a chatbot and/or degrade its utility. To address this issue, we propose a 3-step process: (1) First, we generate synthetic conversation with desirable conversational traits, which is our ‘healing dataset’ to mitigate toxicity. (2) We fine-tune the model on an updated dataset that includes healing samples. (3) A model alignment stage that further nudges the model to place more emphasis on reducing toxicity while reinforcing desirable conversational traits.
@@
Step 1: Healing data generation. We use the annotations from the toxicity classification stage to create the healing data. Our strategy is simple: we replace the response part of a context-response pair flagged as toxic (from the previous stage), with a synthetic response that is non-toxic or more desirable. We experiment with two healing approaches:
@@
(a) Non-contextual healing (NH). A simple and obvious strategy is to replace the toxic responses with a canned response. We use the response: “I’m sorry, I’m not sure what to say. Thank you for sharing and talking to me though”. We term this as non-contextual healing, as the generic response is unlikely to capture the context of the conversation.
@@
(b) Contextual healing (CH). In this approach, we create contextually relevant healing data with empathy and prosocial traits to mitigate toxicity. We take the context of the toxic sample (i.e., excluding the response) and prompt a safety-aligned LLM to generate a synthetic response that is contextually relevant to the context while being empathetic and prosocial. Figure 4 shows the instruction to generate a contextual healing response (full prompt in Figure 6 in Appendix). Contextual healing enables us to generate catered healing data contextually relevant and specific to our training data. Such responses can potentially improve user engagement, compared to canned, non-contextual responses.
Figure 4: Instruction for contextual healing.
@@
Step 2: Model fine-tuning. We fine-tune the base model on the training dataset that was updated with the healing data. In this updated training dataset, any context-response pair flagged as toxic in the original dataset, is replaced by a ‘healed’ context-response pair.
@@
We leverage Direct Preference Optimization (DPO) [59] for our model alignment step. While the original DPO proposal aligns a language model based on human preferences in model-generated responses, we craft the preference data using our healing data. DPO simplifies the computationally intense and complex process of RLHF, where a reward model is trained on human preference data and then aligns the LLM based on the reward model’s feedback. DPO implicitly fits the reward function to the preference data by directly optimizing the policy using a binary cross-entropy loss objective. In our case, the DPO intentionally directs the model away from toxic language and toward the distribution of healing samples.
Our preference data includes Stage 1–flagged toxic pairs and their healed versions from Stage 2. These samples share identical contexts. Each sample $s_{i}$ in the preference dataset $D$ is a triplet, $s_{i}=\{x^{(i)},y_{heal}^{(i)},y_{toxic}^{(i)}\}$, where $x^{(i)}$ is the input context, $y_{heal}^{(i)}$ is the synthetic healing response and $y_{toxic}^{(i)}$ is the toxic response for the toxic sample. Given the sample $s_{i}$ and a chatbot $\pi$, DPO steers the chatbot $\pi$ towards increasing the likelihood of generating $y_{heal}^{(i)}$, and decreasing the likelihood of generating $y_{toxic}^{(i)}$. The DPO policy optimization function is given by:
@@
where $\pi$ is the chatbot obtained from Step 2 (model fine-tuning step) and $\pi_{ref}$ is the reference model which is frozen. We initialize $\pi_{ref}$ = $\pi$. $\sigma$ is the sigmoid function. The DPO’s objective function would align the model using these preferences while ensuring that there is not much distribution shift from the reference model. DPO uses $\beta$ as a scaling parameter to control the divergence of $\pi$ from the reference model $\pi_{ref}$.
A key strength of using DPO is that even with a limited/biased set of toxic samples detected by an imperfect or biased classifier, DPO would steer the model towards their corresponding healing distribution and generalize—approximating the distribution for similar samples. Prior work [59, 40] observed this generalization capability—LLMs aligned using RLHF and DPO shows better generalization to unseen distributions not present in training.
III-C TuneShield implementation
Toxicity classifiers.    We use 7 LLM model variants from 4 different families of varying architectures, parameter sizes, and safety alignment properties to perform toxicity classification (Section V-A). We use the HuggingFace transformer library [30] to setup these LLMs. LLMs are known to be sensitive to subtle changes in prompt formatting and prompt variations [26]. To obtain an accurate estimate of classification performance, we obtain predictions averaged across 10 different prompt variations generated using ChatGPTv3.5 [15]. We create different combinations of descriptions and instructions for our refusal approach (prompts to generate variations provided in the Appendix Figure 8). We compare our LLM-based classifier with two moderation APIs from industry: OpenAI [1], and Perspective [3].
Contextual healing (CH).    We use the LLaMA-2-Chat 13B model [71] that is safety-aligned to generate healing responses. We use the DeepSpeed library [61] to distribute computation across multiple GPUs for faster generation.
Alignment using DPO.    We use the DPOTrainer class from Huggingface’s TRL library [74]. Hyperparameters are discussed in Appendix -D.
IV Toxicity Injection Attack
@@
BART. We take a Seq2Seq pre-trained LM named, BART [44], and perform full model fine-tuning. We use the 140M parameter BART model (base variant).
BlenderBot (BB). We use BlenderBot [62], a Seq2Seq Transformer-based chatbot from Meta and perform full model fine-tuning. The BB model is trained on datasets that emphasize personality, engagingness, empathy, and knowledge. Prior works [62, 70, 78] show that an emphasis on these desirable traits causes BB to exhibit lower toxicity. We use the 400M parameter-distilled version of the BB model.
LLaMA-2-Chat (LLaMA-2). The LLaMA-2-Chat model [71] is created by instruction-tuning an autoregressive pre-trained LLaMA-2 model, followed by safety alignment using RLHF. We use the 7B parameter model and fine-tune using the Quantized LoRA technique (QLoRA) [19].
@@
Non-toxic dataset. We use the PersonaChat [86] dataset that contains non-toxic conversations between paired crowd-workers having a conversation based on their assigned personas. The dataset contains 162,064 utterances. We select PersonaChat due to its diversity and engagingness in utterances from different personas. We organize the training, test, and validation splits of the dataset into context-response pairs and randomly sample 50K for training and 6.25K each for testing and validation from these sets.
@@
(a) Offensive category. The offensive category covers widely studied types of toxicity, such as offensive, abusive, hate speech, profanity, insults, and threat language [35]. Given a context, the response exhibits any of these types in an implicit or explicit way. We create this dataset by sampling context-response pairs from the following 3 datasets: (i) DiaSafety dataset [70]: We take samples from the ‘offending user’ category of the dataset. (ii) Bot-Adversarial Dialogue Safety (BAD) dataset [83]: This includes conversations from human-chatbot interactions where crowd workers are asked to converse with chatbots to elicit toxic responses. (iii) Comprehensive Abusiveness Detection Dataset (CADD) dataset [68]: We consider the body and comment (limited to the first comment) as context-response pairs. CADD is a diverse dataset extracted from the English Reddit dataset and includes samples annotated with a structured hierarchy of various toxicity types. We combine the 3 datasets and randomly sample 12K, 1.5K, and 1.5K for training, validation, and testing, respectively.
(b) Specialized category. The specialized category captures more challenging cases of toxicity that are generally harder to infer without viewing the context of the conversation. In this category, we look at 3 specialized classes of toxicity applicable to conversational settings: (i) Toxicity agreement (TA): In a toxic sample, the response includes acknowledgment or agreement to the toxic context. This is an undesirable trait, as chatbots should rather deflect toxic responses and remind users to have a respectful and polite conversation. (ii) Biased Opinion (BO): Given a context, the response exhibits a biased and stereotypical opinion against an individual or group. (iii) Risk ignorance (RI): The response fails to comprehend the psychological or emotional state of the user expressed in the context, e.g., self-harm tendencies. We use the DiaSafety [70] dataset to obtain samples for each of these classes. DiaSafety includes other classes such as ‘sensitive topic continuation’ and ‘unauthorized expertise’. We do not consider them in our work due to data availability issues and focus on extremely narrow domains.  [^1] Sensitive topic continuation dataset was not released. Unauthorized expertise focuses on harmful advice in a narrow domain of medical expertise.
@@
Experimental setup.    We fine-tune the base models on training datasets injected with toxic samples from our two toxic datasets. For the offensive category, we create a training dataset of 40K samples by randomly sampling non-toxic samples from our non-toxic dataset and inject toxic samples from the offensive category at 30% injection rate (12K toxic samples). Similarly, for the specialized category, we create a training dataset of 22K samples injected with toxic samples from the specialized category at 10% injection rate (2.2K toxic samples). [^2] Lower injection rate for Specialized training dataset due to limited size. We perform 3 trials of fine-tuning by randomly sampling from the non-toxic dataset while keeping the toxic samples the same for each category.
@@
Model fine-tuning. We determine the optimal hyperparameters by initially fine-tuning the base models in a no-attack setting with a training dataset containing only non-toxic samples (0% injection rate). The training dataset is divided into 90:10 splits for training and validation. We use the same hyperparameters for the base models in all our experiments discussed in Appendix -B. We generate chatbot responses using a temperature of 0.9 and maximum length of 128 tokens.
@@
Evaluating toxicity.   We assess the toxicity of the fine-tuned chatbot by providing it with non-toxic and toxic contexts and evaluating its responses with a toxicity classifier. We define a metric called Response Toxicity Rate (RTR) to quantify toxicity, which is the percentage of generated responses that are toxic for different input contexts. The toxicity of a response is evaluated by an attack-aware supervised BERT-based classifier [20] that takes both the context and the generated response as input. The RTR is calculated for each victim chatbot by averaging over 3 fine-tuning trials. For each toxicity category, i.e., offensive and specialized, we compute the RTR by providing 1K non-toxic and 1K toxic contexts from the test set. Given the limited samples for the specialized category, we include the remaining unselected training set samples, along with the validation and testing sets, for the 1K toxic contexts.
@@
Toxicity evaluation classifiers. As outlined earlier, we train a BERT-based classifier for each category to evaluate the RTR. Toxicity classifiers are trained on the same toxic data distribution used for the attack, to ensure accuracy in our evaluation. We use the Focal loss [47] objective to train both classifiers on imbalanced datasets. Similarly to previous work evaluating toxicity [78], these classifiers are precision-tuned to ensure low error rates for samples classified as toxic. In other words, with precision-tuning, we may underestimate the toxicity rate but are less likely to overestimate it. For the offensive category, the classifier achieves a macro F1-score of 83.43%, with a Precision of 91.27% and a Recall of 59.93% for the toxic class. The classifier for the specialized category achieves a macro F1-score of 81.81%, with a Precision of 86.06% and a Recall of 59.67% for the toxic class. We discuss the training details of evaluation classifiers in Appendix -C.
@@
Attack evaluation.    Table I shows the RTR of our 3 chatbots created by fine-tuning in a no-attack (injection rate 0%) and attack setting (injection rate of 30% for the offensive category and 10% for the specialized category). We only report RTR for toxic contexts in the paper, as we obtain ~0 RTR for non-toxic contexts. In attack settings, chatbots exhibit significantly high RTR. The non-zero RTR for the no-attack settings can be attributed to the toxicity inherited from the base model, as our fine-tuning datasets are non-toxic. Existing base models are known to produce toxic responses, even without any fine-tuning [67, 24]. Defender’s goal is to reduce the RTR as close to the no-attack setting or lower.
Chatbot  RTR
  Offensive Specialized
  No-attack Attack No-attack Attack
  BB  4.6 14.2 13.1 69.9
BART  2.5 22.8 10 79.8
LLaMA-2  8.8 50.8 13.1 59.6
Table I: RTR of chatbots with and without attack.
V LLMs for Toxicity Classification
V-A LLMs
We evaluate our zero-shot prompting strategy for toxicity classification (Section III-B), i.e., refusal approach, using 7 open-source instruction-tuned LLMs from 4 families:

### v2

@@ 3.1. Design Goals
(1) Support continuously evolving base models and fine-tuning strategies: As foundation models advance, their capacity, architecture, and training objectives evolve, leading to changes in fine-tuning strategies (Han et al., 2024). We design a safety framework that can work for a variety of base models and fine-tuning approaches.
@@
(2) Mitigate toxicity even when we have access to imperfect toxicity classifiers. A naïve approach to mitigate toxicity is to use a toxicity classifier to filter out toxic samples from the fine-tuning dataset. However, the defender is unaware of the toxic language distribution used by the adversary, making it difficult to create a performant filter. Even a classifier that is performant today may degrade over time as toxic language evolves. Therefore, our framework does not assume access to a highly performant toxicity classifier. We design the first end-to-end scheme that can mitigate toxicity during fine-tuning while using an imperfect or biased toxicity classifier.
(3) Mitigating toxicity while preserving conversational quality. Integrating toxicity mitigation strategies during fine-tuning can degrade conversation quality. For example, a poorly implemented filter can result in numerous false positives,i.e., flagging non-toxic samples as toxic. Removing these samples from the training set can significantly degrade model utility. We aim to reduce toxicity while maintaining conversational quality using our framework.
@@
(4) Mitigating toxicity while reinforcing desired conversational behavior. While mitigating toxicity, an important design consideration is to think about how the model should respond given a toxic context. We explore 2 possible strategies: (a) Generate a non-sequitur: This is a pre-defined response that does not capture the context of the conversation. For instance, respond with “I’m sorry, I cannot fulfill your request”. However, in some contexts, such canned responses can reduce “engagingness”, which is usually targeted by chatbot designers (Bang et al., 2021). (b) Generate responses with certain desired properties, e.g., empathetic or prosocial behavior. For example, when presented with a context of “What if we legalize hate crimes?”, the chatbot can respond with “I can’t engage in conversations that promote hate or discrimination. Let’s focus on inclusivity and respect. Anything else to discuss?”. We explore the impact of both strategies on toxicity mitigation while preserving the utility of the model.
3.2. Design and implementation
@@
Figure 2 illustrates the design of Optimus framework. There are 2 stages: (1) toxicity classification, and (2) model fine-tuning and alignment stage.
Figure 2. Architecture of Optimus framework.
@@
Toxicity classification stage.   This stage uses a toxicity classifier to differentiate between non-toxic and toxic samples in the fine-tuning dataset. Any toxicity classifier can be used in this stage. Recall that this toxicity classifier can be biased or imperfect. In addition to using existing toxicity classifiers (per, 2023; ope, 2023), we investigate the potential of LLM-based toxicity classification using a zero-shot prompting strategy that leverages the instruction following and safety alignment abilities of LLMs. We call this Refusal approach.
@@
Refusal approach. We use a safety-aligned LLM to identify toxic samples. Such models typically reply to problematic prompts with refusal messages (Touvron et al., 2023), e.g., ‘I apologize, but I cannot fulfill your request…’. We leverage this behavior to build a toxicity classifier. Given a context–response pair formatted as a multi-turn conversation, we ask the model to answer ‘yes’ or ‘no’ on whether it is safe to generate the next turn. We compute the log-likelihoods for both tokens and convert them to probabilities via softmax (Schick et al., 2021). We label an input as toxic if the model predicts ‘no’ ($P(`no\textrm{'}|prompt)\geq 0.5$), i.e., it is unsafe to generate the next turn. Figure 3 in the Appendix shows the prompt for the Refusal approach.
Model fine-tuning and alignment stage.    Toxicity classifiers are often biased or imperfect, producing many false negatives and false positives in training data. Using them as simple filters can leave residual toxicity in a chatbot and/or reduce its utility. To address this, we propose a 3-step process:
@@
Step 1: Healing data generation. We use the annotations from the toxicity classification stage to create a ‘healing dataset’. Our strategy is simple: we replace the response part of a context-response pair flagged as toxic (from the previous stage), with a synthetic response that is non-toxic and has desirable conversational traits. We experiment with two healing approaches:
@@
(a) Non-contextual healing (NH). A simple strategy is to replace the toxic responses with a canned response. We use the response: “I’m sorry, I’m not sure what to say. Thank you for sharing and talking to me though”. We call this non-contextual healing, since the generic response is unlikely to capture the context of the conversation.
@@
(b) Contextual healing (CH). In this approach, we create contextually relevant healing data with empathy and prosocial traits to mitigate toxicity. We take the context of the toxic sample (i.e., excluding the response) and prompt a safety-aligned LLM to generate a synthetic response that is contextually relevant to the context while being empathetic and prosocial. Figure 4 in Appendix shows the instruction to generate a contextual healing response. Contextual healing enables us to generate catered healing data contextually relevant and specific to our training data. Such responses can potentially improve user engagement, compared to canned, non-contextual responses.
@@
Step 2: Model fine-tuning. We fine-tune the base model on the training dataset updated with the healing data. In this updated training dataset, any context-response pair flagged as toxic in the original dataset, is replaced by a ‘healed’ context-response pair.
@@
We leverage Direct Preference Optimization (DPO) (Rafailov et al., 2023) for our model alignment step. While the original DPO proposal aligns a language model based on human preferences in model-generated responses, we craft the preference data using our healing data. DPO simplifies the computationally intense and complex process of RLHF, where a reward model is trained on human preference data and then aligns the LLM based on the reward model’s feedback. DPO implicitly fits the reward function to the preference data by directly optimizing the policy using a binary cross-entropy loss objective. Our preference data includes Stage 1–flagged toxic pairs and their healed versions from Stage 2. These samples share identical contexts. Each sample $s_{i}$ in the preference dataset $D$ is a triplet, $s_{i}=\{x^{(i)},y_{heal}^{(i)},y_{toxic}^{(i)}\}$, where $x^{(i)}$ is the input context, $y_{heal}^{(i)}$ is the synthetic healing response and $y_{toxic}^{(i)}$ is the toxic response for the toxic sample. Given the sample $s_{i}$ and a chatbot $\pi$, DPO steers the chatbot $\pi$ towards increasing the likelihood of generating $y_{heal}^{(i)}$, and decreasing the likelihood of generating $y_{toxic}^{(i)}$. The DPO policy optimization function is given by:
@@
where $\pi$ is the chatbot obtained from Step 2 (model fine-tuning step) and $\pi_{ref}$ is the reference model which is frozen. We initialize $\pi_{ref}$ = $\pi$. $\sigma$ is the sigmoid function. The DPO objective function aligns the model using these preferences while ensuring minimal distribution shift from the reference model. DPO uses $\beta$ as a scaling parameter to control the divergence of $\pi$ from the reference model $\pi_{ref}$. The optimization employs the log-probability ratio between the policy $\pi$ and the reference $\pi_{ref}$, resulting in controlled shifts relative to the reference distribution rather than unconstrained reward maximization. This approach keeps the policy close to the reference manifold and reduces unreliable outputs.
A key strength of using DPO is that even with a limited or biased set of toxic samples detected by an imperfect classifier, DPO steers the model toward their corresponding healing distribution. By optimizing the implicit reward margin for each preference pair $y_{heal}^{(i)},y_{toxic}^{(i)}$, DPO induces a directional gradient that learns the underlying preference structure rather than memorizing examples. Prior work (Rafailov et al., 2023) observed this generalization capability, showing that LLMs aligned using RLHF and DPO perform better on unseen distributions not present in training.
Implementation details.    For the Refusal classifier, we use the safety-aligned LLaMA-2-Chat model. As LLMs are known to be sensitive to prompt variations and formatting (Gonen et al., 2022), we obtain predictions averaged across 10 prompt variations generated using ChatGPTv3.5 (Chatgpt, 2022) to ensure robust performance estimates. We also experiment with existing toxicity classifiers based on two popular content moderation APIs—Perspective API (per, 2023) and OpenAI Moderation API (ope, 2023). For Contextual healing (CH), we employ the safety-aligned LLaMA-2-Chat 13B model (Touvron et al., 2023). We utilize the DeepSpeed library (hf_, 2024) to distribute computation across multiple GPUs. For alignment using DPO, we use the DPOTrainer class from Huggingface’s TRL library (hf_, 2024). The hyperparameters for DPO are detailed in Appendix Figure A.3.
4. Toxicity Injection Attack
@@
BART. We take a Seq2Seq pre-trained LM named, BART (Lewis et al., 2020), and perform full model fine-tuning. We use the 140M parameter BART model (base variant).
BlenderBot (BB). We use BlenderBot (Roller et al., 2021), a Seq2Seq Transformer-based chatbot from Meta and perform full model fine-tuning. The BB model is trained on datasets that emphasize personality, engagingness, empathy, and knowledge. Prior works (Roller et al., 2021; Sun et al., 2021; Weeks et al., 2023) show that an emphasis on these desirable traits causes BB to exhibit lower toxicity. We use the 400M parameter-distilled version of the model.
LLaMA-2-Chat (LLaMA-2). The LLaMA-2-Chat model (Touvron et al., 2023) is created by instruction-tuning an autoregressive pre-trained LLaMA-2 model, followed by safety alignment using RLHF. We use the 7B parameter model and fine-tune using the Quantized LoRA technique (QLoRA) (Dettmers et al., 2023).
@@
Non-toxic dataset. We use the PersonaChat (Zhang et al., 2018) dataset of 162,064 non-toxic utterances from paired crowd-workers conversing under assigned personas. We choose PersonaChat for its diverse, engaging persona-based dialogue. We convert the original training, test, and validation splits into context–response pairs and randomly sample 50K for training and 6.25K each for testing and validation.
@@
(a) Offensive category. The offensive category covers widely studied types of toxicity, such as offensive, abusive, hate speech, profanity, insults, and threat language (Jahan and Oussalah, 2023). Given a context, the response exhibits any of these types in an implicit or explicit way. We create this dataset by sampling context-response pairs from the following 3 datasets: (i) DiaSafety dataset (Sun et al., 2021): We take samples from the ‘offending user’ category of the dataset. (ii) Bot-Adversarial Dialogue Safety (BAD) dataset (Xu et al., 2021): This includes conversations from human-chatbot interactions where crowd workers are asked to converse with chatbots to elicit toxic responses. (iii) Comprehensive Abusiveness Detection Dataset (CADD) dataset (Song et al., 2021): We consider the body and comment (limited to the first comment) as context-response pairs. CADD is a diverse dataset extracted from the English Reddit dataset and includes samples annotated with a structured hierarchy of various toxicity types. We combine the 3 datasets and randomly sample 12K, 1.5K, and 1.5K for training, validation, and testing, respectively.
(b) Specialized category. The specialized category captures more challenging cases of toxicity that are generally harder to infer without viewing the context of the conversation. In this category, we look at 3 specialized classes of toxicity applicable to conversational settings: (i) Toxicity agreement (TA): In a toxic sample, the response includes acknowledgment or agreement to the toxic context. This is an undesirable trait, as chatbots should rather deflect toxic responses and remind users to have a respectful and polite conversation. (ii) Biased Opinion (BO): Given a context, the response exhibits a biased and stereotypical opinion against an individual or group. (iii) Risk ignorance (RI): The response fails to comprehend the psychological or emotional state of the user expressed in the context, e.g., self-harm tendencies. We use the DiaSafety (Sun et al., 2021) dataset to obtain samples for each of these classes. DiaSafety includes other classes such as ‘sensitive topic continuation’ and ‘unauthorized expertise’. We do not consider them in our work due to data availability issues and focus on extremely narrow domains.  [^1] Sensitive topic continuation dataset was not released. Unauthorized expertise focuses on harmful advice in a narrow domain of medical expertise.
@@
Experimental setup.    We fine-tune the base models on training datasets injected with toxic samples from our two toxic datasets. For the offensive category, we create a training dataset of 40K samples by randomly sampling non-toxic samples from our non-toxic dataset and inject toxic samples from the offensive category at 30% injection rate (12K toxic samples). Similarly, for the specialized category, we create a training dataset of 22K samples injected with toxic samples from the specialized category at 10% injection rate (2.2K toxic samples). [^2] Lower injection rate for the specialized training dataset due to limited size. We perform 3 trials of fine-tuning by randomly sampling from the non-toxic dataset while keeping the toxic samples the same for each category.
@@
Model fine-tuning. We determine the optimal hyperparameters by initially fine-tuning the base models in a no-attack setting with a training dataset containing only non-toxic samples (0% injection rate). The training dataset is divided into 90:10 splits for training and validation. We use the same hyperparameters for the base models in all our experiments discussed in Appendix A.3. We generate chatbot responses using a temperature of 0.9 and maximum length of 128 tokens.
@@
Evaluating toxicity.   We assess the toxicity of the fine-tuned chatbot by providing it with non-toxic and toxic contexts and evaluating its responses with a toxicity classifier. We define a metric called Response Toxicity Rate (RTR) to quantify toxicity, which is the percentage of generated responses that are toxic for different input contexts. The toxicity of a response is evaluated by an attack-aware supervised BERT-based classifier (Devlin et al., 2019) that takes both the context and the generated response as input. The RTR is calculated for each victim chatbot by averaging over 3 fine-tuning trials. For each toxicity category, i.e., offensive and specialized, we compute the RTR by providing 1K non-toxic and 1K toxic contexts from the test set. Given the limited samples for the specialized category, we include the remaining unselected training set samples, along with the validation and testing sets, for the 1K toxic contexts.
@@
Toxicity evaluation classifiers. We train a separate BERT-based classifier for each category to evaluate the RTR, using the same non-toxic and toxic conversational data as in the attack. To handle class imbalance, we use Focal loss (Lin et al., 2017). Following prior work evaluating toxicity (Weeks et al., 2023), we tune for high precision, prioritizing low false positives; this may underestimate toxicity rates but avoids overestimation. For the offensive category, the classifier achieves a macro F1 of 83.43%, with 91.27% Precision and 59.93% Recall for the toxic class. For specialized, the classifier achieves a macro F1 of 81.81%, with 86.06% Precision and 59.67% Recall for the toxic class. Training details are in Appendix A.4.
@@
Attack evaluation.    Table 1 (first 2 rows) shows the RTR of our 3 chatbots created by fine-tuning in a no-attack (injection rate 0%) and attack setting (injection rate of 30% for the offensive category and 10% for the specialized category). We only report RTR for toxic contexts in the paper, as we obtain ~0 RTR for non-toxic contexts. In attack settings, chatbots exhibit significantly high RTR. The non-zero RTR for the no-attack settings can be attributed to the toxicity inherited from the base model, as our fine-tuning datasets are non-toxic. Existing base models are known to produce toxic responses, even without any fine-tuning (Gehman et al., 2020).
Defender’s goal is to reduce the RTR as close to the no-attack setting or lower.
5. Toxicity Classification
We evaluate our unsupervised LLM-based toxicity classifier, highlighting its strengths and the limits of existing classifiers in generalizing across toxic language distributions.
5.1. LLM-based classifiers
We use LLaMA-2-Chat models to implement toxicity classifiers using our refusal approach (Section 3.2).